If I have 2.9 L of gas at a pressure of 5.0 atm and a temperature of 50.0 °C, what will be the new temperature if I decrease the volume of the gas to 2.4 L and decrease the pressure to 3.0 atm(round to correct sig figs)


1
Expert's answer
2021-08-22T07:29:48-0400

pV=nRT

n = pV / RT

R = 0.082 057 46(14) L atm K−1 mol−1

n = 5.0 x 2.9 / (0.082 x (50 + 273)) = 0.5 mol


T = pV / nR

T2 = 3.0 x 2.4 / (0.5 x 0.082) = 175.6 K = -97.4°C